News
Apr
27
Wakfield ACI Cloud Platform Issues – 27th April 2017 – 11:45 BST
Posted by Glenn Akester on 27 April 2017 11:54 AM

Node4’s Network Management Systems have alerted to potential issues with our Wakefield ACI cloud platform,

Due to the resilient nature of our cloud infrastructure customers should only see minor issues as failover events occur.

Engineers are currently investigating and updates will be provided as they become available.

 


 

Updates

27/04/17 12:08 BST – Software instabilities with one of our SDN nodes caused a reload, we will continue to investigate root cause and raise with vendor technical assistance.


Read more »



Apr
26
N4Protect Firewall Issues – 26th April 2017 – 14:54 BST
Posted by Glenn Akester on 26 April 2017 03:08 PM

Node4 Network Management Systems have alerted to potential issues with N4Protect firewall services,

Engineers are currently investigating and updates will be provided as they become available.


Updates

26/04/17 15:20 BST – Node4 engineers believe the issues with N4Protect firewall services are now resolved, any customers still experiencing issues are urged to raise with Node4’s Service Desk. Node4 apologise for any inconvenience caused and will continue analysing root cause, with an RFO to be provided once completed.


Read more »



Apr
4
SIPLink Issue – DC3 (Wakefield / Leeds) cluster
Posted by Eddie Shortt on 04 April 2017 10:48 AM

Our monitoring has alerted us to an issue with the DC3 (Wakefield / Leeds) SIPLink cluster, our engineers are currently investigating and will have an update shortly.

 

The cluster based in DC2 (Derby data centre) is completely unaffected and continues to operate as normal.

 

Update – 11:02 – Our engineers have now identified and resolved the issue. The cluster appears to be stable again and is routing calls correctly, we will continue to investigate the cause and act accordingly.


Read more »



Mar
30
Node4 3G/4G Network Issues – 30/03/17 13:30
Posted by Simon Bell on 30 March 2017 01:47 PM

We are currently investigating issue’s with our 3G/4G connectivity.

We Apologise for any inconvenience this is causing.

Regards

Node4 Technical Support

 

Update — 9:20am 31/03/2017:

The issue is only affect a few customer sites with SIM offline. We are still investigating this with 3G/4G service provider.

 

This issue was caused by EE’s outage and already resolved. If you still experiencing any problem, please contact Node4 Support on 0845 123 2229.


Read more »



Feb
15
Network Issues – 15/02/17
Posted by Alessandro Rizzo on 15 February 2017 09:47 PM

At 21:04 Node4 NOC has reported 10Gbps circuit down within our network, traffic has re-routed around resilient connections.

We are currently in talks with supplier to ascertain the reason for outage.

21:46 – Supplier is aware of the outage we are currently awaiting an update from the carrier.

23:54 – We have had reports from the supplier that there has been a fibre break. They are working to resolve this more updates follow.

01:10 – Our provider has updated us stated that engineers are due on site to the fibre break ETA 03:00am we have no ETA to resolve once they are on site more updates to follow.

03:40 – Our provider have advised that the engineers are now onsite at the fibre break and are working to fix this issue, further updates to follow.

05:20 – Engineers are currently still working to fix the fiber breaks, further updates to follow.

07:35 – Engineers are still working to fix this fiber break. Our provider is unable to provide an ETA for resolution at this time. Further updates to follow.

08:30 – Engineers are still working to fix this fiber break. No ETA as yet. We will update again in an hour.

09:30 – Provider incident Management team and techs are currently on a conference call, we expect an update imminently.

10:26 – Update from provider: – Engineers now need to gain access to a junction chamber located on the corner of a country road in order to continue working on this fiber break. This will require a traffic management team. Traffic management have been contacted and are on route. further updates to follow.

11:26 – Update from provider: –

All prep work is done, joint identified, specialist fibre engineers/OTDR’s etc on site,  still awaiting access to chamber via traffic management.

13:05 – Having chased our provider, we are informed that the status of this fault remains unchanged. Engineers on site are waiting on an emergency permit to work from the local authority. We are actively pursuing an ETA.

14:24 – Update from our providers: Permit and traffic will be in place from 15:00. Engineers are ready to work as soon as access is provided.

We will seek a further update and push for an ETA on fix at 17:00

17:30 – Our provider states that engineers are still on site, continuing to work to splice the fibres. Unfortunately no ETA is yet available on fix.

18:00 – Update from our provider :

The cable has multiple cuts beyond the chamber due to vandalism.

This means we will have to pull in approx. 650 metres of cable and re-splice both ends.

Current ETA for completion is 8 hours ETA approx. 02:00.

22:30 – Update from provider, Cable will be laid and prepped at 00:00. Splicing and testing will commence between 02:00 and 03:00.

02:20 – Update from provider, Cable preparation has been completed as scheduled. Splicing and testing is now taking place. Further updates to follow.

04:00 – Our provider has advised us that all 96 fibers have been successfully spliced, our monitoring is reflecting this as all circuits are up and passing traffic. Engineers are currently carrying our final checks and a further update will follow.

04:30 – Our provider has advised us that all testing has been completed and services have been resumed to normal operation. If you have any queries regarding this outage or have any outstanding issues please get in touch with our service desk at 08451232229.


Read more »



Feb
7
Network Instabilities – 07/02/17 14:10
Posted by Chris Pagel on 07 February 2017 02:23 PM

We are currently investigating network instabilities affecting our London POP – further updates to follow.

 

## Update 15:00

Incident details: Instability began at 14:10 – resolved by 14:32

Services Affected: MPLS services terminating on an MPLS P/PE router in Telehouse North POP

While we are continuing to investigate this issue we have traced the event to a network loop in an external suppliers network (connected to the MPLS router).


Read more »